# Efficient Methods for Non-stationary Online Learning

### v2

@@ 2.1 A Brief Review of Dynamic Regret Minimization
Zhang et al. (2018a) propose a two-layer online algorithm called Ader with an $\mathcal{O}(\sqrt{T(1+P_{T})})$ dynamic regret, which is proved to be minimax optimal for convex functions. Ader maintains a group of base-learners, each performing online gradient descent (OGD) (Zinkevich, 2003) with a customized step size specified by the pool =$\mathcal{H}=\{\eta_{1},\ldots,\eta_{N}\}$, and then uses a meta-algorithm to combine them all. Denoted by $\mathcal{B}_{1},\ldots,\mathcal{B}_{N}$ the $N$ base-learners. For each $i\in[N]$, the base-learner $\mathcal{B}_{i}$ updates by
@@ 2.2 The Reduction Mechanism for Reducing Projection Complexity
where $S_{\mathcal{X}}(\mathbf{y})=\inf_{\mathbf{x}\in\mathcal{X}}\lVert\mathbf{y}-\mathbf{x}\rVert_{2}$ is the distance function to $\mathcal{X}$ and $\mathbf{v}_{t}=\nicefrac{{(\mathbf{y}_{t}-\mathbf{x}_{t})}}{{\lVert\mathbf{y}_{t}-\mathbf{x}_{t}\rVert_{2}}}$ is the vector indicating the projection direction.
@@
0:  step size pool =$\mathcal{H}=\{\eta_{1},\ldots,\eta_{N}\}$, learning rate of meta-algorithm $\varepsilon_{t}$ (or simply a fixed one $\varepsilon_{t}=\varepsilon$).
@@
Set the step size pool as =$\mathcal{H}=\big{\{}\eta_{i}=2^{i-1}(D/G)\sqrt{5/(2T)}\mid i\in[N]\big{\}}$ with $N=\lceil 2^{-1}\log_{2}(1+2T/5)\rceil+1$ and the learning rate as $\varepsilon=\sqrt{(\ln N)/(1+G^{2}D^{2}T)}$. Under Assumptions 1 and 2, our algorithm requires one projection onto $\mathcal{X}$ per round and enjoys
@@
Set the step size pool as =$\mathcal{H}=\big{\{}\eta_{i}=2^{i-1}\sqrt{5D^{2}/(1+8LGDT)}\mid i\in[N]\big{\}}$ with $N=\lceil 2^{-1}\log_{2}((5D^{2}+2D^{2}T)(1+8LGDT)/(5D^{2}))\rceil+1$ and the learning rate of the meta-algorithm as $\varepsilon_{t}=\sqrt{(\ln N)/(1+D^{2}\sum_{s=1}^{t-1}\|\nabla g_{s}(\mathbf{y}_{s})\|_{2}^{2})}$. Under Assumptions 12, and 3, our algorithm requires one projection onto $\mathcal{X}$ per round and enjoys the following dynamic regret:
@@
where $F^{\mathbf{u}}_{T}=\sum_{t=1}^{T}f_{t}(\mathbf{u}_{t})$ is the cumulative loss of the comparators.
@@ 3.1 A Brief Review of Adaptive Regret Minimization
However, there are also some pitfalls. SACS also suffers from an $\mathcal{O}(\log T)$ projection complexity in the worst case due to its two-layer structure. Further, it can be observed that SACS only attains an $\mathcal{O}(\sqrt{\lvert I\rvert\log{\lvert I\rvert}\log T})$ bound in the worst case, which exhibits an $\mathcal{O}(\sqrt{\log\lvert I\rvert})$ gap compared with the best known result of $\mathcal{O}(\sqrt{\lvert I\rvert\log T})$ (Jun et al., 2017).
@@ 3.2 Efficient Algorithms for Adaptive Regret
Adaptive regret examines every contiguous interval $I\subseteq[T]$, which demands a rapid adaptation to potential environment changes. A natural way to construct the scheduling is to initiate a base-learner at each round and enable her to make predictions till the end of the whole time horizon (Hazan and Seshadhri, 2007). While this approach can effectively maintain many base-learners to handle the non-stationarity, it is computationally expensive due to maintaining $\mathcal{O}(t)$ base-learners at round $t$. For enhanced efficiency, alternative scheduling is proposed to set the length of each learner’s active time in a geometric manner (Hazan and Seshadhri, 2007; Daniely et al., 2015). To elucidate the design, we introduce two concepts: the unit interval and the marker. The unit intervals partition the time horizon $[T]$, of which the adaptive algorithm chooses geometric many to construct the active time as illustrated later. The markers denote the starting and ending time stamps of the unit intervals. Formally, the $i$-th unit interval is represented by $[s_{i},s_{i+1}-1]$ with the time stamps ${s_{1},\dots,s_{M}}$ determining the intervals (referred to as markers). Notice that, these unit intervals are disjoint and consecutive. The second and third rows of Figure 1 provide an illustrative example of unit intervals and markers.
@@
where $m=i\cdot 2^{k},k\in\{0\}\cup\mathbb{N}$ and $k$ is the largest number made the factorization. The setting indicates that the base-learner initialized at time $s_{m}$ remains active across $12^{k}-$ 1 unit intervals, leading to an exponentially number of unit intervals. This scheduling diversifies multiple base-learners to capture the non-stationarity across different time durations, and meanwhile ensures that at most $\mathcal{O}(\log T)$ base-learners are maintained per round.
@@
10:        Initialize a new base-learner whose active span is $[s_{n\cdot 2^{k}},s_{(n+1)\cdot 2^{k}}-1]$ where $m=n\cdot 2^{k},k\in\{0\}\cup\mathbb{N}$ and $k$ is the largest number made the factorization.
@@
SACS uses the AdaNormalHedge (Luo and Schapire, 2015) as the meta-algorithm, however, this is not suitable for our proposal. To ensure one projection per iteration, we cannot use multiple function values, i.e., $\{g_{t}(\mathbf{y}_{t,i})\}_{i=1}^{\lvert A_{t}\rvert}$, for meta-algorithm to evaluate the loss. Instead, we can only use the linearized loss value, namely, $\{\langle\nabla g_{t}(\mathbf{y}_{t}),\mathbf{y}_{t,i}\rangle\}_{i=1}^{\lvert A_{t}\rvert}$ in the weight update of meta-algorithm. The small-loss regret bound in the meta-algorithm of SACS crucially relies on the original function values, which is unfortunately inaccessible in our case. Technically, when fed with linearized loss, it is hard to establish a squared gradient-norm bound and then convert it to the small loss due to the first-order regret bound of AdaNormalHedge. Based on this crucial technical observation, we propose to use the Adapt-ML-Prod algorithm (Gaillard et al., 2014) as the meta-algorithm in our method. The key advantage is that it enjoys a second-order regret and also supports the sleeping expert setup. Adapt-ML-Prod maintains multiple learning rates $\bm{\eta}_{t+1}\in[0,1]^{\lvert A_{t+1}\rvert}$ and an intermediate weight vector $\mathbf{w}_{t+1}\in\mathbb{R}^{\lvert A_{t+1}\rvert}$, which are updated by the following rule. For any active base-learner $i\in A_{t+1}$,
@@
^$\widehat{\ell}_{t}=\langle\nabla g_{t}(\mathbf{y}_{t}),\mathbf{y}_{t}\rangle/(2GD),\mbox{ and }\ell_{t,i}=\langle\nabla g_{t}(\mathbf{y}_{t}),\mathbf{y}_{t,i}\rangle/(2GD).$ (13)
@@ 4.1 A Brief Review of Interval Dynamic Regret Minimization
Cutkosky (2020) achieves an $\widetilde{\mathcal{O}}(\sqrt{\lvert I\rvert(1+P_{I})})$ interval dynamic regret through a fundamentally different framework based on parameter-free online learning, where $\widetilde{\mathcal{O}}(\cdot)$ hides the polynomial dependence on $\log T$ factors. In fact, Cutkosky (2020) derives a gradient-norm bound (scaling with $\sum_{t\in I}\lVert\nabla f_{t}(\mathbf{x}_{t})\rVert^{2}$) for interval dynamic regret, which can be easily converted into a small-loss bound (scaling with $\sum_{t\in I}f_{t}(\mathbf{u}_{t})$) when the online functions are non-negative and smooth. However, we note that, even in this case, their small-loss bound is not favorable enough, as the regret guarantee is not fully problem-dependent — it still exhibits polynomial dependence on $\log T$ terms.
@@ 4.2 Efficient Algorithms for Interval Dynamic Regret
Regarding the small-loss type bound, the only relevant result in the literature is the gradient-norm bound presented by Cutkosky (2020), which can be further converted into a small-loss bound of order $\widetilde{\mathcal{O}}(\sqrt{(F_{I}^{\mathbf{u}}+P_{I})(1+P_{I})})$ when the online functions are non-negative and smooth, where $F_{I}^{\mathbf{u}}=\sum_{t\in I}f_{t}(\mathbf{u}_{t})$ and $P_{I}=\sum_{t\in I}\lVert\mathbf{u}_{t}-\mathbf{u}_{t-1}\rVert_{2}$. However, it is only a partially problem-dependent bound, as the $\widetilde{\mathcal{O}}(\cdot)$ notation still hides polynomial dependence on $\log T$ factors.
@@
0:  step size pool =$\mathcal{H}=\{\eta_{1},\ldots,\eta_{N}\}$, learning rate setting $\varepsilon_{i,t}$ for the $i$-th base-learner, and the threshold generating function $\mathcal{G}(\cdot):\mathbb{N}\mapsto\mathbb{R}_{+}$.
@@
10:        Initialize a new base-learner according to Eq. (17) and Eq. (18), whose active span is $[s_{n\cdot 2^{k}},s_{(n+1)\cdot 2^{k}}-1]$ where $m=n\cdot 2^{k},k\in\{0\}\cup\mathbb{N}$ and $k$ is the largest number made the factorization.
@@
[…]
We employ the prefix “base:” to signify that these formulas are updated as part of the base-algorithm utilized by Algorithm 3. The local decision $\mathbf{y}_{t+1,i,j}$ is returned from the $j$-th base:base-learner in the base-algorithm, $N$ denotes the number of base:base-learners, $i_{\text{start}}$ denotes the starting time stamp for the $i$-th base-learner and $p_{t+1,i,j}$ denotes its corresponding combination weight. Note that these base-level updates are conducted within the domain $\mathcal{Y}$, where the algorithm can benefit from a rapid projection. Moreover, it is also worth mentioning the choice of step size $\eta_{j}$. Ideally, it should be scaled with the length of the active interval $\lvert I\rvert$ of the associated base-learner, nevertheless, we actually set it based on $T$ in Theorem 5 due to the problem-dependent covering intervals. Technically, since the active time of each base-learner is determined on the fly, we can only set potentially over-estimated learning rates to ensure that the dynamic regret is guaranteed whenever the base-learner is deactivated. As stated in Lemma 8, such a configuration equips base-learners with an anytime dynamic regret bound.
Bringing all components together, in the following theorem, we establish the small-loss interval dynamic regret for Algorithm 3.
@@
Under Assumptions 13, employing the step size pool for base:base-algorithm (defined at Eq. (17)) as =$\mathcal{H}=\{\eta_{j}=2^{j-1}\sqrt{5D^{2}/(1+8LGDT)}\mid j\in[N]\}$ with $N=\lceil 2^{-1}\log_{2}((5D^{2}+2D^{2}T)(1+8LGDT)/(5D^{2}))\rceil+1$, setting the meta-algorithm’s learning rate of the $i$-th base-learner (defined at Eq. (18)) as $\varepsilon_{i,t}=\sqrt{(\ln N)/(1+D^{2}\sum_{s=i_{\text{start}}}^{t-1}\|\nabla g_{s}(\mathbf{y}_{s})\|_{2}^{2})}$ with $i_{\text{start}}$ denoting its starting time stamp, setting the threshold generating function
$\begin{split}\mathcal{G}(m)={}&7L\left(12D\sqrt{\ln(1+2m)}+4D\mu_{T}+6D\sqrt{\ln N}\right)^{2}\\ &{}+54GD\ln(1+2m)+18GD\mu_{T}+\frac{3(6+G^{2}D^{2})\sqrt{\ln N}}{2}+(630L+23)D^{2}+9,\end{split}$ (19)
@@
$\mathcal{O}\left(\sqrt{(F^{\mathbf{u}}_{I}+P_{I})\cdot\big{(}P_{I}+\log\big{(}\min_{\mathbf{u}_{1:T}}\{F^{\mathbf{u}}_{T}+P_{T}\}\big{)}\big{)}\cdot\log\big{(}\min_{\mathbf{u}_{r:s}}\{F^{\mathbf{u}}_{I}+P_{I}\}\big{)}}\right),$ (20)
@@
Algorithm 3 actually enjoys the same gradient-norm bound as Cutkosky (2020, Theorem 7) (in fact even stronger by logarithmic factors in $T$) under Assumptions 1 and 2 (without requiring the smoothness assumption). More specifically, our analysis is conducted in terms of $\lVert\nabla g_{t}(\mathbf{y}_{t})\rVert_{2}$ (see Eq. (65) and Eq. (66)), which can then be related to $\lVert\nabla f_{t}(\mathbf{x}_{t})\rVert_{2}$ by Lemma 1. Following the same convention of Cutkosky (2020), who uses the $\widetilde{\mathcal{O}}(\cdot)$-notation to hide the logarithmic dependence in $T$, Algorithm 3 can obtain an $\widetilde{\mathcal{O}}(\sqrt{(G_{I}+P_{I})(P_{I}+1)})$ interval dynamic regret for any interval $I=[r,s]\subseteq[T]$ without any modification on the learning rates, where $P_{I}=\sum_{t=r+1}^{s}\|\mathbf{u}_{t}-\mathbf{u}_{t-1}\|_{2}$ and $G_{I}=\sum_{t=r}^{s}\|\nabla f_{t}(\mathbf{x}_{t})\|_{2}^{2}$.
@@
The interval dynamic regret bound obtained by Cutkosky (2020) essentially implies a small-loss bound $\widetilde{\mathcal{O}}(\sqrt{(F^{\mathbf{u}}_{I}+P_{I})(1+P_{I})})$ when assuming the smoothness of loss functions. However, the $\widetilde{\mathcal{O}}(\cdot)$ notation hides problem-independent factors such as $\log T$. In contrast, our result in Theorem 5 is fully problem-dependent, replacing the quantity such as $\log T$ by $\log(\min_{\mathbf{u}_{1:T}}\{F^{\mathbf{u}}_{T}+P_{T}\})$, through our techniques of constructing problem-dependent covering intervals and corresponding threshold designs. These problem-dependent quantities can be much smaller under nice environments, and it remains unclear how to adapt our techniques to the method of Cutkosky (2020) due to a different framework.

### v3

@@ 2.1 A Brief Review of Dynamic Regret Minimization
Zhang et al. (2018a) propose a two-layer online algorithm called Ader with an $\mathcal{O}(\sqrt{T(1+P_{T})})$ dynamic regret, which is proved to be minimax optimal for convex functions. Ader maintains a group of base-learners, each performing online gradient descent (OGD) (Zinkevich, 2003) with a customized step size specified by the pool $\mathcal{H}=\{\eta_{1},\ldots,\eta_{N}\}$, and then uses a meta-algorithm to combine them all. Denoted by $\mathcal{B}_{1},\ldots,\mathcal{B}_{N}$ the $N$ base-learners. For each $i\in[N]$, the base-learner $\mathcal{B}_{i}$ updates by
@@ 2.2 The Reduction Mechanism for Reducing Projection Complexity
where $S_{\mathcal{X}}(\mathbf{y})=\inf_{\mathbf{x}\in\mathcal{X}}\lVert\mathbf{y}-\mathbf{x}\rVert_{2}$ is the distance function to $\mathcal{X}$ and $\mathbf{v}_{t}=(\mathbf{y}_{t}-\mathbf{x}_{t})/\lVert\mathbf{y}_{t}-\mathbf{x}_{t}\rVert_{2}$ is the vector indicating the projection direction.
@@
0: step size pool $\mathcal{H}=\{\eta_{1},\ldots,\eta_{N}\}$, learning rate of meta-algorithm $\varepsilon_{t}$ (or simply a fixed one $\varepsilon_{t}=\varepsilon$).
@@
Set the step size pool as $\mathcal{H}=\big{\{}\eta_{i}=2^{i-1}(D/G)\sqrt{5/(2T)}\mid i\in[N]\big{\}}$ with $N=\lceil 2^{-1}\log_{2}(1+2T/5)\rceil+1$ and the learning rate as $\varepsilon=\sqrt{(\ln N)/(1+G^{2}D^{2}T)}$. Under Assumptions 1 and 2, our algorithm requires one projection onto $\mathcal{X}$ per round and enjoys
@@
Set the step size pool as $\mathcal{H}=\big{\{}\eta_{i}=2^{i-1}\sqrt{5D^{2}/(1+8LGDT)}\mid i\in[N]\big{\}}$ with $N=\lceil 2^{-1}\log_{2}((5D^{2}+2D^{2}T)(1+8LGDT)/(5D^{2}))\rceil+1$ and the learning rate of the meta-algorithm as $\varepsilon_{t}=\sqrt{(\ln N)/(1+D^{2}\sum_{s=1}^{t-1}\|\nabla g_{s}(\mathbf{y}_{s})\|_{2}^{2})}$. Under Assumptions 12, and 3, our algorithm requires one projection onto $\mathcal{X}$ per round and enjoys the following dynamic regret:
@@
where $F^{\mathbf{u}}_{T}=\sum_{t=1}^{T}f_{t}(\mathbf{u}_{t})$ is the cumulative loss of the dynamic comparators.
@@ 3.1 A Brief Review of Adaptive Regret Minimization
[…]
@@ 3.2 Efficient Algorithms for Adaptive Regret
Adaptive regret examines every contiguous interval $I\subseteq[T]$, which demands a rapid adaptation to potential environment changes. A natural way to construct the scheduling is to initiate a base-learner at each round and enable her to make predictions till the end of the whole time horizon (Hazan and Seshadhri, 2007). While this approach can effectively maintain many base-learners to handle the non-stationarity, it is computationally expensive due to maintaining $\mathcal{O}(t)$ base-learners at round $t$. For enhanced efficiency, alternative scheduling is proposed to set the length of each learner’s active time in a geometric manner (Hazan and Seshadhri, 2007; Daniely et al., 2015). To elucidate the design, we introduce two concepts: the unit interval and the marker. The unit intervals partition the time horizon $[T]$, of which the adaptive algorithm chooses geometric many to construct the active time as illustrated later. The markers denote the starting and ending time stamps of the unit intervals. Formally, the $i$-th unit interval is represented by $s_{i},s_{i+1}-1]$ with the time stamps $s_{1},\dots,s_{M}$} determining the intervals (referred to as markers). Notice that, these unit intervals are disjoint and consecutive. The second and third rows of Figure 1 provide an illustrative example of unit intervals and markers.
@@
where $m=i\cdot 2^{k},k\in\{0\}\cup\mathbb{N}$ and $k$ is the largest number made the factorization. The setting indicates that the base-learner initialized at time $s_{m}$ remains active across $2^{k}-1$ unit intervals, leading to an exponentially number of unit intervals. This scheduling diversifies multiple base-learners to capture the non-stationarity across different time durations, and meanwhile ensures that at most $\mathcal{O}(\log T)$ base-learners are maintained per round.
@@
10:   Initialize a new base-learner whose active span is $s_{n\cdot 2^{k}},s_{(n+1)\cdot 2^{k}}-1]$ where $m=n\cdot 2^{k},k\in\{0\}\cup\mathbb{N}$ and $k$ is the largest number made the factorization.
@@
SACS uses AdaNormalHedge (Luo and Schapire, 2015) as its meta-algorithm, but this choice is not suitable for us. To ensure only one projection per iteration, we cannot rely on multiple function values $\{g_{t}(\mathbf{y}_{t,i})\}_{i=1}^{\lvert A_{t}\rvert}$ for the meta-algorithm to evaluate the loss. Instead, we can only use the linearized loss value, namely, $\{\langle\nabla g_{t}(\mathbf{y}_{t}),\mathbf{y}_{t,i}\rangle\}_{i=1}^{\lvert A_{t}\rvert}$ in the meta update. The small-loss bound in SACS crucially depends on access to the original function values, which are unavailable here. Technically, with only linearized losses, it is difficult to establish a squared gradient-norm bound and then convert it to the small-loss bound due to the first-order regret bound of AdaNormalHedge. This key observation motivates us to instead use the Adapt-ML-Prod algorithm (Gaillard et al., 2014) as our meta-algorithm. The main advantage is that it enjoys a second-order regret while also supporting the sleeping expert setup. Adapt-ML-Prod maintains multiple learning rates $\bm{\eta}_{t+1}\in[0,1]^{\lvert A_{t+1}\rvert}$ and an intermediate weight vector $\mathbf{w}_{t+1}\in\mathbb{R}^{\lvert A_{t+1}\rvert}$, , updated as follows: for any active base-learner $i\in A_{t+1}$,
@@
$\widehat{\ell}_{t}=\langle\nabla g_{t}(\mathbf{y}_{t}),\mathbf{y}_{t}\rangle/(2GD),\mbox{ and }\ell_{t,i}=\langle\nabla g_{t}(\mathbf{y}_{t}),\mathbf{y}_{t,i}\rangle/(2GD).$ (13)
@@ 4.1 A Brief Review of Interval Dynamic Regret Minimization
[…]
@@ 4.2 Efficient Algorithms for Interval Dynamic Regret
[…]
@@
0: step size pool $\mathcal{H}=\{\eta_{1},\ldots,\eta_{N}\}$, learning rate setting $\varepsilon_{i,t}$ for the $i$-th base-learner, and the threshold generating function $\mathcal{G}(\cdot):\mathbb{N}\mapsto\mathbb{R}_{+}$.
@@
10:   Initialize a new base-learner according to Eq. (17) and Eq. (18), whose active span is $s_{n\cdot 2^{k}},s_{(n+1)\cdot 2^{k}}-1]$ where $m=n\cdot 2^{k},k\in\{0\}\cup\mathbb{N}$ and $k$ is the largest number made the factorization.
@@
$\texttt{base:base-level:}\quad\mathbf{y}_{t+1,i,j}$ $=\Pi_{\mathcal{Y}}\left[\mathbf{y}_{t+1,i,j}-\eta_{j}\nabla g_{t}(\mathbf{y}_{t})\right],j\in[N],$ (17)
$\texttt{base:meta-level:}\quad p^{\prime}_{t+1,i,j}$ $\propto\exp\left(-\varepsilon_{i,t+1}\sum_{s=i_{\text{start}}}^{t}\langle\nabla g_{s}(\mathbf{y}_{s}),\mathbf{y}_{s,i,j}\rangle\right),j\in[N],$ (18)
We use the prefix “base:” to indicate updates performed within the base-algorithm utilized by Algorithm 3. The local decision $\mathbf{y}_{t+1,i,j}$ is returned from the $j$-th base:base-learner in the base-algorithm, $N$ denotes the number of base:base-learners, $i_{\text{start}}$ denotes the starting time stamp for the $i$-th base-learner and $p_{t+1,i,j}$ denotes its corresponding combination weight. Note that these base-level updates are conducted within the domain $\mathcal{Y}$, where the algorithm can benefit from a rapid projection. Moreover, it is also worth mentioning the choice of step size $\eta_{j}$. Ideally, it should be scaled with the length of the active interval $\lvert I\rvert$ of the associated base-learner, nevertheless, we actually set it based on $T$ in Theorem 5 due to the problem-dependent covering intervals. Technically, since the active time of each base-learner is determined on the fly, we can only set potentially over-estimated learning rates to ensure that the dynamic regret is guaranteed whenever the base-learner is deactivated. As shown in Lemma 8, this setup ensures that each base-learner enjoys an anytime dynamic regret bound.
With these components in place, the following theorem establishes the small-loss interval dynamic regret of Algorithm 3.
@@
Under Assumptions 13, employing the step size pool for base:base-algorithm (defined at Eq. (17)) as $\mathcal{H}=\{\eta_{j}=2^{j-1}\sqrt{5D^{2}/(1+8LGDT)}\mid j\in[N]\}$ with $N=\lceil 2^{-1}\log_{2}((5D^{2}+2D^{2}T)(1+8LGDT)/(5D^{2}))\rceil+1$, setting the meta-algorithm’s learning rate of the $i$-th base-learner (defined at Eq. (18)) as $\varepsilon_{i,t}=\sqrt{(\ln N)/(1+D^{2}\sum_{s=i_{\text{start}}}^{t-1}\|\nabla g_{s}(\mathbf{y}_{s})\|_{2}^{2})}$ with $i_{\text{start}}$ denoting its starting time stamp, setting the threshold generating function
$\begin{split}\mathcal{G}(m){}&=7L\left(12D\sqrt{\ln(1+2m)}+4D\mu_{T}+6D\sqrt{\ln N}\right)^{2}\\ &{}+54GD\ln(1+2m)+18GD\mu_{T}+\frac{3(6+G^{2}D^{2})\sqrt{\ln N}}{2}+(630L+23)D^{2}+9,\end{split}$ (19)
@@
$\mathcal{O}\bigg{(}\sqrt{(F^{\mathbf{u}}_{I}+P_{I})\cdot\big{(}P_{I}+\log\big{(}\min_{\mathbf{u}_{1:T}}\{F^{\mathbf{u}}_{T}+P_{T}\}\big{)}\big{)}\cdot\log\big{(}\min_{\mathbf{u}_{r:s}}\{F^{\mathbf{u}}_{I}+P_{I}\}\big{)}}\bigg{)},$ (20)
@@
[…]
@@
[…]